x^2+y^2-12x-24y+36=0
Answer by stanbon(75887)

x^2+y^2-12x-24y+36=0
---
Complete the square on the x and on the y terms.
---
x^2-12x + ? + y^2 -24y + ? = -36
---
x^2 - 12x + 6^2 + y^2 - 24y + 12^2 = -36 + 6^2 + 12^2
(x-6)^2 + (y-12)^2 = 12^2
(x-6)^/144 + (y-12)^2/144 = 1
====================================
Cheers,
Stan H.